OpenStudy (anonymous):

10 over 2- square root of 5 given this radical expression, rationalize the denominator .

OpenStudy (anonymous):

sorry, it's -(20+10sqrt5)

OpenStudy (anonymous):

wouldnt it be 20+10 sqrt 5/ -3 or -20-10 sqrt 5

OpenStudy (anonymous):

yes, if you distribute the negative sign that is what you would get

OpenStudy (anonymous):

so -20-10 sqrt 5

OpenStudy (anonymous):

to rationalize stuff like this, just multiply both the numberator and the denominator by the conjugate of what term with the sqrt in it. Yes, that is right.

OpenStudy (anonymous):

For this one 3 sqrt of -4 lies between two consecutive integers what are the integers.... I don't know how to start this? Would you happen to know.

OpenStudy (anonymous):

I know it has to do with imaginary numbers, but I haven't done that since beginning of high school. I'm sure if you post it as a question someone else will be able to help you out.

OpenStudy (anonymous):

Yeah, I haven't done this in so LONG. But I will post it, thanks for your help!

OpenStudy (anonymous):

Anytime! Good luck

OpenStudy (radar):

\[10\over2-\sqrt{5}\]Is this the problem?

OpenStudy (anonymous):

Yes.

OpenStudy (radar):

O.K multiply top and bottom by\[2+\sqrt{5}\]this will give you \[10(2+\sqrt{5})\over -1\] or \[-20-10\sqrt{5}\]Actually there is no imaginary component to this it is a irrational number.

OpenStudy (anonymous):

the imaginary comment was about a different problem.

OpenStudy (anonymous):

Oh we figured this one out, we were trying to figure out another equation that is as follows 3 square root of -4 lies between two consecutive integers what are the two consecutive integers.

OpenStudy (radar):

Oh very well, good luck with this, and complex numbers with imaginary component are quite useful in the field of electronics.

OpenStudy (anonymous):

\[\sqrt[3]{-4}\] the A is a 3...That is what the equation looks like.

OpenStudy (anonymous):

sqrt(-4) is 2i

OpenStudy (anonymous):

I know the square root of 4 is 2. That is not my question... My question is what two consecutive integers lie between the equation 3 square root of -4....

OpenStudy (radar):

\[\sqrt[3]{-4}\] Is also not imaginary

OpenStudy (radar):

The solution is a negative value real and irrational.

OpenStudy (radar):

-1.587401052